# Viscoelastic Worthington jets & droplets produced by bursting bubbles

### v1

@@ 5 Conclusion and outlook
Analysis across the $Ec$-$De$ phase space reveals markedly different behaviors in two limiting cases. For polymeric liquids with permanent memory ($De\to\infty$), transitions occur at fixed $Ec$, independently of $De$. The jetting transition $Ec_{j}$ is independent of $Oh_{s}$, while the dropping transition $Ec_{d}$ exhibits strong $Oh_{s}$ dependence. Remarkably, this infinite $De$ asymptote extends down to $De\approx 1$ due to the finite process timescale. Conversely, for polymeric liquids with poor memory ($De\to 0$), both transitions occur at constant polymeric Ohnesorge number $Oh_{p}=Ec\times De$, indicating that addition of polymers introduces an excess viscous stress in this limit. These transitions are independent of $Oh_{s}$. Using a slender jet approach (Driessen et al., 2013; Gordillo et al., 2020; Eggers & Fontelos, 2015), we provide further insights into these transitions, examining the competition between elastic stresses and inertial flow focusing that governs jet formation and droplet ejection. This analysis helps to explain the observed scaling behaviors and transition criteria.
@@
Our findings have important implications for understanding and controlling bubble bursting in viscoelastic fluids, with relevance to biological processes (Walls et al., 2017), such as airborne disease transmission (Bourouiba, 2021), and industrial applications, such as inkjet printing (Lohse, 2022). The results highlight how polymer additives can dramatically alter spray formation, potentially enhancing or suppressing aerosol production depending on the fluid properties (Kant et al., 2023). This work also opens several avenues for future research. Further investigation is needed into the universal behavior near the infinite $De$ asymptote, including the development of closed-form $De$ expansions for free surface flows (Sen et al., 2021; França et al., 2024; Sen et al., 2024; Balasubramanian et al., 2024; Boyko et al., 2024; Hinch et al., 2024). The mechanism underlying the $Oh_{s}$ sensitivity of transition $Ec$ values at high $De$ requires further clarification. Additionally, extending our analysis to nonlinear viscoelastic models would provide valuable insights into the role of shear-thinning behavior and finite extensibility on bursting bubbles, addressing limitations of the current model (Zinelis et al., 2023; McKinley & Sridhar, 2002; Snoeijer et al., 2020). This approach would allow quantification of discrepancies between experiments and simulations, often attributed to inherent issues with the Oldroyd-B model, thereby enhancing our understanding of viscoelastic jets (Gaillard et al., 2024). Indeed, the numerical method developed here, freely available at Sanjay & Dixit (2024), provides a generalized framework readily adaptable to any model within the Oldroyd-B family of upper convective derivative models (Snoeijer et al., 2020). Furthermore, as higher Bond numbers are observed in many scenarios (Walls et al., 2015; Ghabache et al., 2014; Deike et al., 2018; Krishnan et al., 2017), exploring their combined effect with viscoelasticity on the overall dynamics would provide valuable insights into such experiments (Rodríguez-Díaz et al., 2023). Lastly, studying interactions of multiple bubbles (Singh & Das, 2019) at the liquid-gas free surface will provide further insights into pathogen transport.
@@
In conclusion, this study provides a comprehensive characterization of bubble bursting in viscoelastic media, interpreting the interplay between elastic, viscous, and capillary forces by moving in the $Oh_{s}$-$Ec$-$De$ phase space. The insights gained advance our fundamental understanding of free-surface flows in complex fluids and pave the way for more precise control of spray formation in various applications.
@@
Acknowledgments. We would like to thank Jacco Snoeijer, Maziyar Jalaal, Uddalok Sen, Andrea Prosperetti, and Vincent Bertin for the discussions.

### v2

@@ 5 Conclusion and outlook
Analysis across the $Ec$-$De$ phase space reveals markedly different behaviors in two limiting cases. For polymeric liquids with permanent memory ($De\to\infty$), transitions occur at fixed $Ec$, independently of $De$. The jetting transition $Ec_{j}$ is independent of $Oh_{s}$, while the dropping transition $Ec_{d}$ exhibits strong $Oh_{s}$ dependence. Remarkably, this infinite $De$ asymptote extends down to $De\approx 1$, where the polymer relaxation time becomes comparable to the process timescale. Below this, for $De\sim\mathcal{O}(0.1)$, we observe a transition in scaling behavior, consistent with the Weissenberg number criterion $Wi\equiv De\sqrt{We_{\text{jet}}}\sim\mathcal{O}(1)$, where $We_{\text{jet}}$ is the jet Weber number (equation (24)) that remains approximately constant due to negligible elastic effects during the initial shear flow (Blanco-Rodríguez & Gordillo, 2021). Conversely, for polymeric liquids with poor memory ($De\to 0$), both transitions occur at constant polymeric Ohnesorge number $Oh_{p}=Ec\times De$, indicating that the addition of polymers introduces an excess viscous stress in this limit. These transitions are independent of $Oh_{s}$. Using a slender jet approach (Driessen et al., 2013; Gordillo et al., 2020; Eggers & Fontelos, 2015), we provide further insights into these transitions, examining the competition between elastic stresses and inertial flow focusing that governs jet formation and droplet ejection. This analysis helps to explain the observed scaling behaviors and transition criteria.
@@
Our findings have important implications for understanding and controlling bubble bursting in viscoelastic fluids, with relevance to biological processes (Walls et al., 2017), such as airborne disease transmission (Bourouiba, 2021), and industrial applications, such as inkjet printing (Lohse, 2022). The results highlight how polymer additives can dramatically alter spray formation, with intermediate values of $Ec$ and $De$ leading to smaller and faster droplets, whereas high values of $Ec$ and $De$ suppress droplet formation entirely (Kant et al., 2023). This work also opens several avenues for future research. Further investigation is needed into the universal behavior near the infinite $De$ asymptote, including the development of closed-form $De$ expansions for free surface flows (Sen et al., 2021; França et al., 2024; Sen et al., 2024; Boyko et al., 2024; Hinch et al., 2024). The mechanism underlying the $Oh_{s}$ sensitivity of transition $Ec$ values at high $De$ requires further clarification. Additionally, extending our analysis to nonlinear viscoelastic models would provide valuable insights into the role of shear-thinning behavior and finite extensibility on bursting bubbles, addressing limitations of the current model (Zinelis et al., 2024; McKinley & Sridhar, 2002; Snoeijer et al., 2020). This approach would allow quantification of discrepancies between experiments and simulations, often attributed to inherent issues with the Oldroyd-B model, thereby enhancing our understanding of viscoelastic jets (Gaillard et al., 2024b). Indeed, the numerical method developed here, freely available at Sanjay & Dixit (2024), provides a generalized framework readily adaptable to any model within the Oldroyd-B family of upper convective derivative models (Snoeijer et al., 2020). Furthermore, as higher Bond numbers are observed in many scenarios (Walls et al., 2015; Ghabache et al., 2014; Deike et al., 2018; Krishnan et al., 2017), exploring their combined effect with viscoelasticity on the overall dynamics would provide valuable insights into such experiments (Rodríguez-Díaz et al., 2023). Indeed, a critical assumption of this work is the initial condition and its history, particularly for bubbles at liquid-gas interfaces in viscoelastic or elastoviscoplastic media. Our current work assumes the bubble has resided at the interface for a duration far exceeding the polymeric medium’s relaxation time, ensuring elastic stresses have fully relaxed before bursting. This idealized scenario provides a well-defined starting point but may not fully capture experimental conditions (Cheny & Walters, 1996; Deoclecio et al., 2023). Lastly, studying interactions of multiple bubbles (Singh & Das, 2019) at the liquid-gas free surface will provide further insights into pathogen transport.
@@
In conclusion, this study investigates and characterizes bubble bursting in viscoelastic media, interpreting the interplay between elastic, viscous, and capillary forces by moving in the $Oh_{s}$-$Ec$-$De$ phase space. As a starting point, we employed the Oldroyd-B constitutive model. While this choice elucidates the basic interplay of elasticity, viscosity, and capillarity, it does not capture shear-thinning effects or finite extensibility of polymer chains. Therefore, the predicted droplet sizes, jet thinning dynamics, and ultimate filament breakup must be interpreted with caution. More complex viscoelastic models (e.g., Giesekus, FENE-P) that incorporate finite extensibility and nonlinearities will likely alter certain details of our findings. Hence, our results should be viewed as a conceptual road map rather than definitive predictions. An essential extension of our study involves the experimental validation of the numerical results. Controlled laboratory studies using polymer solutions with known rheological properties are needed to assess the accuracy of the Oldroyd-B model in this parameter regime (also see appendix B). Such comparisons will help determine where the simplified assumptions fail and guide refinements, including the use of more realistic constitutive equations.
Despite these caveats, our study offers a foundation for understanding how viscoelasticity can either suppress or enhance droplet formation during bubble bursting. We hope this work will inspire future experiments and numerical explorations using more advanced rheological models, ultimately leading to a more complete and quantitative picture of viscoelastic bubble bursting across different application domains.
@@
Acknowledgments. We would like to thank Andrea Prosperetti, Gareth McKinley, Jacco Snoeijer, Maziyar Jalaal, Uddalok Sen, and Vincent Bertin for discussions.
@@ Appendix A The Newtonian limit of bursting bubble dynamics
Appendix B A note on the range of control parameters considered in this work
In this appendix, we tabulate and compare the range of dimensionless parameters explored in this work with those available in the literature on viscoelastic effects in bubble bursting. Tables B 1 and B 2 summarize the physical properties and corresponding dimensionless numbers from three representative experimental studies.
$c$ $R$ $\eta_{s}$ $\lambda$ $\eta_{p}$ $G$
[…]
Cheny & Walters (1996) [0, 100] 7.5, 19 300 N/A [0, 18] N/A
Rodríguez-Díaz et al. (2023) [0, 350] 1 1 [0, 500] [0, 0.5] [0, 1]
Cabalgante-Corralesa et al. (2024) [0, 100] 0.93 0.89 [0, 700] [0, 2] [0, 1]
[…]
$Oh_{s}$ $De$ $Ec$ $Oh_{p}$ $Bo$
This work [10-3, $10^{0}$] [0, $\infty$) [0, 103] [0, $\infty$) $10^{-3}$
Balasubramanian et al. (2024) [$10^{-3},10^{-2}$] [0, $10^{2}$] [0, $10$] [$10^{-3},10^{-2}$] $10^{-3}$
Cheny & Walters (1996) $10^{-1}$ N/A N/A [0, $10^{-2}$] [$10,10^{2}$]
Rodríguez-Díaz et al. (2023) $10^{-3}$ [0, $10^{-1}$] [0, $10^{-2}$] [0, $10^{-3}$] $10^{-1}$
Cabalgante-Corralesa et al. (2024) $10^{-3}$ [0, $2\times 10^{-1}$] [0, $10^{-2}$] [0, $10^{-2}$] $10^{-1}$
Table B 2: Representative values of dimensionless numbers in this work as compared to those from previous studies. For experimental studies, the dimensionless parameters are calculated using the properties in table0B 1. For Balasubramanian et al. (2024), we have only considered the limiting cases of zero yield-stress. We note that while experiments are naturally limited in their accessible parameter ranges, our numerical study explores a broader range to establish comprehensive scaling laws and regime transitions.
Table B 1 presents key physical parameters including polymer concentration ($c$), bubble radius ($R$), solvent viscosity ($\eta_{s}$), polymer relaxation time ($\lambda$), polymer contribution to viscosity ($\eta_{p}$), and elastic modulus ($G$). The corresponding dimensionless numbers are shown in Table B 2, where we compare our parameter space with both experimental and computational studies from the literature. Our work systematically explores a significantly broader range of these parameters compared to experimental studies, which are often constrained by practical limitations in achievable polymer concentrations and relaxation times. This comprehensive coverage allows us to identify universal scaling laws and regime transitions that may be challenging to observe experimentally.
The ranges explored in our numerical study suggest several promising directions for future experimental investigations. For instance, while moving in the $De$-$Ec$ parameter space, experiments could probe the robustness of our predicted transitions and scaling laws. Experimental studies would not only validate our computational findings but could also reveal additional physical mechanisms not captured by the Oldroyd-B model. We anticipate that trying new polymers and advances in characterization techniques (Gaillard et al., 2024a) will continue to expand the experimentally accessible parameter space, enabling increasingly detailed comparisons between simulations and experiment.
Appendix C Grid sensitivity tests
This appendix assesses the grid independence of our numerical results by examining two important metrics: (i) the predicted droplet size and (ii) the regime transitions. Ensuring grid convergence is crucial, especially if interface ruptures due to finite grid resolution in our numerical code (Lohse & Villermaux, 2020; Chirco et al., 2022; Kant et al., 2023).
Figure C 1(a) shows the relative error in predicted droplet size as a function of the number of grid points per initial bubble radius $R_{0}/\Delta$, where $\Delta$ is the minimum grid size. We focus on $De\to\infty$ as this case is particularly demanding, featuring slender filaments due to viscoelastic stresses. The error is calculated relative to the finest resolution ($R_{0}/\Delta=2048$). The data exhibit approximately first-order convergence, indicated by the dashed line scaling as $(R_{0}/\Delta)^{-1}$. For our standard resolution of $R_{0}/\Delta=512$, the relative error is approximately 6%, decreasing to about 3% at $R_{0}/\Delta=1024$.
While droplet size convergence demonstrates improved numerical accuracy with increasing resolution, the determination of regime transitions between different flow behaviors provides an even more stringent test. These transitions are highly sensitive to the details of jet breakup. Figure C 1(b) displays the dimensionless elastocapillary number $Ec$ at the transition boundary for different grid resolutions. We find that for $(R_{0}/\Delta)\geq 1024$, the transition curves do not change, confirming that the scaling behaviors previously identified – namely $Ec_{d}\sim De^{-1}$ for $De\ll 1$ and $Ec_{d}\sim De^{0}$ for $De\gg 1$ – are robustly reproduced across all grid resolutions tested.
Figure C 1: (a) Relative error in predicted droplet size versus the number of grid points per bubble radius, $R_{0}/\Delta$, at $De\to\infty$. The dashed line indicates a scaling of $(R_{0}/\Delta)^{-1}$, demonstrating approximately first-order convergence. (b) Dependence of the critical elastocapillary number $Ec_{d}$ at the dropping transition on the Deborah number $De$ for different grid resolutions ($R_{0}/\Delta=256,512,1024,2048$). The scaling behaviors $Ec_{d}\sim De^{-1}$ as $De\to 0$ and $Ec_{d}\sim De^{0}$ as $De\to\infty$ remain unchanged beyond $R_{0}/\Delta=1024$.
Appendix D Deviation from the Newtonian asymptote
Figure D 1: Comparison of (a) maximum jet length $L_{\text{max}}/R_{0}$ against $Ec$ at various $De$ at fixed representative cases of $Oh_{s}=0.05$ and (b) first droplet size $r_{d}/R_{0}$ against $Ec$ at various $De$ fixed at $Oh_{s}=0.001$. The horizontal lines indicate the Newtonian reference values (obtained at $Ec=0$). At small $Ec$, both $L_{\text{max}}$ and $r_{d}$ coincide with their Newtonian counterparts, demonstrating negligible viscoelastic influence. As $Ec$ increases beyond critical values, significant deviations from the Newtonian limits emerge, with the degree of departure depending on $De$. These results quantify the onset and magnitude of elastic effects relative to the Newtonian baseline, providing a clear framework for interpreting viscoelastic modifications to bursting bubble dynamics.
In the main text, we showed that for small elastocapillary numbers $Ec$, the droplet size $r_{d}$ and jet length $L_{\text{max}}$ closely match those of the Newtonian case at arbitary $Oh_{s}$ (solvent Ohnesorge number) and $De$ (Deborah number). Only when $Ec$ approaches or exceeds critical values do we observe significant departures from the Newtonian reference.
Figure D 1 quantifies these deviations by comparing both the maximum jet length $L_{\text{max}}$ (Figure D 1a) and the first droplet size $r_{d}$ (Figure D 1b) against $Ec$ at various $De$, in the limit of $Oh_{s}\ll 1$. The symbols represent numerical results for the viscoelastic system, while the horizontal lines mark the corresponding Newtonian asymptotes (i.e., $r_{d}$ and $L_{\text{max}}$ values obtained at $Ec=0$). For small $Ec$, both $r_{d}$ and $L_{\text{max}}$ are invariant, indicating that viscoelastic stresses are negligible in this range. As $Ec$ increases and approaches the critical thresholds identified in § 4, deviations emerge, ultimately leading to suppressed jetting or droplet formation.
Notably, the critical $Ec$ value at which $r_{d}$ and $L_{\text{max}}$ deviate from their Newtonian counterparts depends on $De$. For high $De$, even a moderate increase in $Ec$ can trigger significant changes, reflecting the persistent elastic memory in the fluid. In contrast, for $De\ll 1$, where the polymeric stresses relax rapidly, larger $Ec$ values are necessary to produce noticeable departures from Newtonian behavior. Similarly, the Newtonian limit is readily recovered by reducing either $Ec$ or $De$ to zero.
These results highlight that any interpretation of viscoelastic bubble-bursting dynamics should be framed with reference to the Newtonian baselines (either $De=0$ or $Ec=0$). By systematically mapping out these deviations, one can pinpoint the onset of non-Newtonian behavior and interpret observed jetting or droplet formation regimes as outcomes of either weak or strong elastic effects, all benchmarked against the Newtonian scenario.